OpenStudy (anonymous):

How do I show the inverse of 2 functions graphically?

OpenStudy (anonymous):

f(x)=x+6 g(x)=x-6

OpenStudy (agent0smith):

You can find the inverse of each and graph them each... or graph the original function, and reflect it over the line y=x

OpenStudy (anonymous):

OH I SEE NOW!

OpenStudy (anonymous):

what do i do for this part though? show that they are inverse by verifying that (fog)(x)=x and (gof)(x)=x

OpenStudy (agent0smith):

fog(x) means find f(g(x)) which means... f(x-6). gof(x) means g(f(x)) so find g(x+6).

OpenStudy (anonymous):

okay thank you